OpenStudy (anonymous):

Evaluate the summation of 20 times 0.5 to the n minus 1 power, from n equals 3 to 12..

OpenStudy (jdoe0001):

\(\bf {\huge \Sigma}_{{\color{brown}{ n}}=3}^{12}\quad a_1\cdot r^{{\color{brown}{ n}}-1}\implies a_1\left(\cfrac{1-r^{\color{brown}{ n}}}{1-r}\right)\)

OpenStudy (anonymous):

can you explain further my level of confusion is insane @jdoe0001

OpenStudy (jdoe0001):

well... \(\bf a_1 = 20\ and \ r=0.5\) \(\bf a_1=\textit{1st term of sequence}\quad r=\textit{common ratio or "multiplier"}\)

OpenStudy (anonymous):

okay so what is the first term of sequence??

OpenStudy (jdoe0001):

so... what you have is pretty much... includes the 1st term and the common ratio thus \(\bf {\huge \Sigma}_{{\color{brown}{ n}}=3}^{{\color{blue}{ 12}}}\quad 20\cdot 0.5^{{\color{brown}{ n}}-1}\implies (20\cdot 0.5^{{\color{brown}{ 3}}-1})\left(\cfrac{1-0.5^{\color{brown}{ {\color{blue}{ 12}}}}}{1-0.5}\right)\)

OpenStudy (jdoe0001):

the 1st term... .in the sequence is 20 HOWEVER you're asked to start adding from the n = 3, or 3rd term thus I used \(\bf (20\cdot 0.5^{{\color{brown}{ 3}}-1})\)

OpenStudy (anonymous):

okay so what is next

OpenStudy (jdoe0001):

get the summation :)

OpenStudy (anonymous):

how ???/

OpenStudy (jdoe0001):

\(\bf {\huge \Sigma}_{{\color{brown}{ n}}=3}^{{\color{blue}{ 12}}}\quad 20\cdot 0.5^{{\color{brown}{ n}}-1} \\ \quad \\ \implies \Large (20\cdot 0.5^{{\color{brown}{ 3}}-1})\left(\cfrac{1-0.5^{\color{brown}{ {\color{blue}{ 12}}}}}{1-0.5}\right)\)

OpenStudy (anonymous):

9.99 19.95 29.98 39.96 so how do i get to one of these

OpenStudy (jdoe0001):

well.... what did you get for the one above?

OpenStudy (jdoe0001):

if you need a calculator https://www.desmos.com/calculator

OpenStudy (anonymous):

thats the thing idk how to solve it.. im bad at math and my mom cant help me right now... so i have 0 clue what im doing with this

OpenStudy (jdoe0001):

hmm well.. the idea is that you have to calculate the value so....

OpenStudy (jdoe0001):

say for example.... what does \(\bf (20\cdot 0.5^{{\color{brown}{ 3}}-1})\) give you?

OpenStudy (anonymous):

1.5

OpenStudy (jdoe0001):

hm..... may want to recheck it

OpenStudy (anonymous):

5

OpenStudy (jdoe0001):

so.... 5 yes try now just the numerator of the fracton :)

OpenStudy (anonymous):

.9997

OpenStudy (jdoe0001):

yeap and the bottom will be 1-0.5 which is just 0.5 so just do a 5 * 0.9997 / 0.5 :)

OpenStudy (anonymous):

9.997 OMG THANK YOU SO MUCHHHHH

OpenStudy (jdoe0001):

yw